OpenStudy (anonymous):

Suppose that a^2 varies inversely with b^3 . If a=7 when b=3 , find the value of a^2 when b=6.

OpenStudy (dumbcow):

\[a^{2} = \frac{k}{b^{3}}\] \[k = a^{2}*b^{3} = 49*27\] \[a^{2} = \frac{49*27}{6^{3}}\]